OpenStudy (anonymous):

What is the slope and y-intercept of the equation 3(y − 2) + 6(x + 1) − 2 = 0? A. slope = -2, y-intercept = 2/3 B. slope = 2, y-intercept = 2/3 C. slope = -2, y-intercept = -2/3 D. slope = 2, y-intercept = -2/3 E. slope = -2, y-intercept = 3/2 PLEASE HELP

Nnesha (nnesha):

first distribute bracket by 3 and other by 6

OpenStudy (anonymous):

3y-6+6x+6-2=0

Nnesha (nnesha):

yes now combine like terms

OpenStudy (anonymous):

3y+6x-2=0

Nnesha (nnesha):

yes now solve for y first add 2 both side

OpenStudy (anonymous):

so 3y+6x=2

Nnesha (nnesha):

subtract 6x both side

Nnesha (nnesha):

then divide both side by 3

OpenStudy (anonymous):

why subtract 6?

Nnesha (nnesha):

you have to solve for y bec slope intercept equation is y =mx +b

Nnesha (nnesha):

not just 6 subtract 6x

OpenStudy (anonymous):

so is y=2/3?

Nnesha (nnesha):

nope

Nnesha (nnesha):

what you get after subtracting 6x both side

OpenStudy (anonymous):

-6x+3y=2?

Nnesha (nnesha):

nope you have to subtract 6x both side that's how you shuld subtract